# SuperMat: Physically Consistent PBR Material Estimation at Interactive Rates

### v3

@@ Abstract
Figure 1: We present an efficient method for high-quality material decomposition in both 2D and 3D. Left: Albedo and RM predictions of SuperMat on rendered images under unknown illumination. The first row shows inputs with uneven lighting, the second row displays albedo predictions, and the third row presents RM predictions. Middle: Rendering results of a 3D object with PBR materials obtained by our method under novel lighting. Right: Our method is efficient and produces high-quality results.
@@ 1 Introduction
Decomposing physically-based materials from images into their constituent properties (e.g., albedo, metallic, and roughness maps) remains a central challenge in computer vision and graphics. This task is essential for creating high-fidelity 3D assets across numerous applications, from film production and game development to virtual and augmented reality pipelines. While recent advances have shown promising results [5, 41, 23], the problem remains challenging due to its ill-posed nature, particularly given unknown lighting conditions and complex interactions between different material properties. Our work addresses two key objectives: 1) efficient decomposition of high-fidelity material maps from 2D images, and 2) extension of this capability to 3D objects, enabling accurate material estimation across complex geometries.
@@
Previous works in this domain generally fall into two main categories: optimization-based and data-driven approaches. 1) Optimization-based methods [2, 43, 44, 15, 27, 16, 11, 21, 42, 14, 6] employ differentiable rendering and neural representations to reconstruct 3D objects with material parameters by minimizing the difference between rendered and target images. While these methods achieve reasonable view consistency, they often compromise material quality, producing artifacts such as blurred boundaries and physically implausible parameters, particularly in metallic and roughness properties. 2) Data-driven approaches [36, 47, 20], particularly recent ones [41, 5, 23], leverage large diffusion models and reframe material decomposition as a conditional image generation task. However, these methods tend to focus solely on individual material properties, overlooking the interactions between different types of materials and combined impact on rendered appearances. Even minor inaccuracies in individual material estimates can propagate through the rendering pipeline, leading to substantial visual artifacts. Moreover, current methods face considerable computational overhead - optimization-based approaches require costly geometric optimization, while diffusion-based methods demand multiple denoising steps (30-50) during inference, causing a critical bottleneck for real-time applications.
@@
To address these challenges, we propose a deterministic one-step framework that achieves both efficient and high-quality material decomposition. First, inspired by HyperHuman [24], we introduce structural expert branches within the UNet architecture. These specialized branches enable concurrent processing of multiple material properties while maintaining domain-specific expertise, allowing our model to decompose albedo, metallic, and roughness maps simultaneously with minimal computational overhead. This design significantly improves consistency between different material properties while avoiding the linear cost scaling typically required when estimating multiple material attributes. Second, we address the computational inefficiency of traditional diffusion models by reformulating the DDIM sampling process [12, 37]. Through careful analysis and modification of the scheduler, we achieve single-step inference without compromising quality, dramatically reducing inference time. Most importantly, this single-step formulation enables end-to-end training of the diffusion model, allowing us to compute the perceptual and re-render loss on final predicted material textures, which enhances shadow removal and improves the joint performance of decomposed materials under varying lighting conditions. These advances culminate in SuperMat, a highly efficient image space material decomposition model that estimates high-fidelity albedo, metallic, and roughness maps at millisecond-scale speeds.
@@
For practical 3D applications, we further extend SuperMat’s capabilities to handle 3D objects. We design a UV refinement network that processes SuperMat’s per-view predictions in UV space, addressing the challenge of maintaining consistency across different viewpoints while preserving the efficiency of our 2D approach. Integrated with SuperMat, this pipeline processes a textured 3D model in approximately 3 seconds, producing high-quality material maps that maintain consistency across arbitrary viewpoints and lighting conditions.
@@
  • With the help of structural expert branches and single-step inference, we successfully achieve end-to-end training with perceptual loss and re-render loss on diffusion models in the material decomposition process, significantly enhancing both material prediction quality and rendering fidelity.
  • We propose an efficient PBR material decomposition method for 3D models. Combining SuperMat and our UV refinement network, this pipeline generates high-quality PBR materials for 3D objects from baked textures in approximately 3 seconds.
@@ 2 Related Works
Material Decomposition. Recognizing the importance of material information in 3D assets, researchers have increasingly focused on material decomposition for objects and scenes. Several inverse rendering approaches [42, 14, 6] aim to reconstruct material-rich 3D representations from single or multi-view images under fixed lighting models. Although assumptions involving extra constraints can simplify the problem and yield valuable insights, they also limit the practical applicability of these methods. Other works [2, 43, 44, 15, 27, 16, 11, 21] focus on estimating both material properties and unknown illumination. As available information decreases, researchers have found that in addition to inverse control over material generation through rendered results, directly acquiring image space material priors is feasible. This shift has inspired a new wave of material decomposition methods [36, 47, 20, 41, 5, 23, 25, 39] targeting 2D images captured or rendered under unknown lighting conditions. However, challenges remain, as limitations in training data, network architecture, and training methods continue to leave room for substantial improvements in decomposition quality.
Diffusion Models. Recently, researchers have applied diffusion models to a range of tasks [29, 4, 35, 30, 45], demonstrating the effectiveness of casting outcome prediction as a conditional image generation process. By leveraging prior knowledge from large pre-trained models, recent studies [13, 34, 18, 9, 46, 10, 38, 5, 23] have achieved performance improvements over previous methods in tasks like depth, normal, and material estimation. However, incorporating large models introduces efficiency challenges, particularly high memory requirements and the time-intensive, multi-step denoising process needed for inference. Inspired by insights from recent work [24, 12, 22], our approach addresses these issues by significantly reducing memory and inference time, improving the feasibility of applying these models in real-world applications.
Generation in UV Space. UV maps are advantageous for fully encapsulating the appearance information of 3D models, leading several works to explore 3D object manipulation directly within the UV space. By representing the texture of an object as a learnable image, these methods [32, 3, 40] enable texture generation, inpainting, and editing, providing new avenues for 3D tasks and allowing 2D image-space techniques and insights to be applied to 3D appearances. Building on this approach, recent methods [1, 26] have applied diffusion models to refine UV maps with promising results. In this spirit, to extend SuperMat’s capability to 3D, we fine-tune a UV-space refinement diffusion model specifically.
@@ 3 Method
In this paper, we propose SuperMat, an image space material decomposition model fine-tuned from Stable Diffusion that efficiently generates high-quality albedo, metallic, and roughness maps simultaneously. Given a rendered or captured image of an object under an unknown lighting condition from a specific viewpoint, SuperMat generates the corresponding material maps for the model from that same viewpoint. As shown in Figure 2, SuperMat is a 2D image-to-image one-step model with structural expert branches that simultaneously generates both albedo and RM (Roughness, Metallic) maps, supporting single-step inference. Compared to previous diffusion-based methods, this model produces superior results in less time and with reduced memory usage. Our single-step formulation allows direct generation of material maps during training, in contrast to prior diffusion methods that predict noise ($\epsilon$) or velocity (v) terms. This enables end-to-end optimization with both perceptual loss on the generated maps and re-render loss from novel lighting conditions, leading to improved material quality and rendering fidelity.
While SuperMat achieves efficient single-image material decomposition, extending this capability to 3D objects presents additional challenges. Instead of relying on optimization-based approaches like score distillation sampling (SDS) that are computationally intensive and prone to mode collapse, we propose an efficient two-stage pipeline for material estimation on 3D objects. Given a 3D model with only RGB texture, we first apply SuperMat to decompose materials from multiple viewpoints. We then introduce a UV refinement network that consolidates these view-dependent predictions into complete, high-quality UV texture maps through single-step inference. As shown in Figure 3, this pipeline enables PBR material estimation for 3D objects in approximately 3 seconds per model.
Next, we first introduce the rendering model we use, followed by a detailed description of all the key components.
3.1 Physically Based Rendering (PBR) Model
We use the Cook-Torrance Bidirectional Reflectance Distribution Function (BRDF) [7] based on microfacet theory to define the materials and establish the rendering model. Following [4], for a point with coordinates $p\in\mathbb{R}^{3}$, albedo $a\in\mathbb{R}^{3}$, metallic $m\in\mathbb{R}$, roughness $r\in\mathbb{R}$, and normal variation $n\in\mathbb{R}^{3}$, the PBR result $L$ observed from viewpoint $c\in\mathbb{R}^{3}$ is given by:
@@ 3.2 SuperMat - Image Space Material Decomposition
Previous works [41, 5, 23] have demonstrated that casting material estimation as a conditional latent diffusion process, leveraging knowledge from large pre-trained diffusion models like Stable Diffusion [33], is effective. However, for our task that requires simultaneous generation of multiple material maps, three key issues remain: 1) High model overhead: Fine-tuning Stable Diffusion with its original architecture can only estimate a single material type. Even when combining metallic and roughness into an RM map, treating it as a single material still requires two diffusion models, one for albedo and another for RM, to meet the rendering model’s requirements. This results in doubled training time and increased inference space overhead. 2) Slow inference speed: With the widely used DDIM scheduler, diffusion models require 30 to 50 denoising steps to produce satisfactory results, which can be inefficient and time-consuming. 3) Incomplete decomposition effect: Due to the high computational cost of backpropagation in multi-step inference, prior diffusion-based methods trained on predicted noise from single-step denoising rather than directly supervising the fully denoised material output. This limitation restricts the model’s ability to specialize in the decomposition task and impedes the application of more task-specific techniques.
@@
Structural Expert Branches. Our solution to the first problem is to introduce a structural expert branch architecture [24]. Previous methods are limited to using a single diffusion model to estimate one material type. Although training separate models for different material maps is an intuitive approach, it introduces issues such as implementation complexity, high overhead, and inconsistencies between materials. Considering that, despite their different representations, materials fundamentally reflect an understanding of object features, we incorporate structural expert branches into the UNet model, allowing a single diffusion model to predict multiple material maps. Specifically, we design two expert branches, one for predicting albedo, and the other for predicting RM. For each expert branch, we replicate the UNet’s last UpBlock, and conv_out, while the rest of the UNet architecture remains shared between the branches. The modified model first extracts and transforms encoded image features through the shared modules. These features are then separated into 2 copies by duplicating before the last UpBlock, with each branch ultimately predicting results for its respective domain. The expert branches specialize in material-specific features, while the shared modules focus on commonalities. In this way, we only add a small number of parameters, achieving the same performance as previous methods that would otherwise require twice the scale.
Single-Step Inference and End-to-End Training. To address the second problem, we analyze the DDIM scheduler of diffusion models. Previous works have highlighted a flaw in the implementation of the DDIM scheduler [12, 22], which limits the potential of single-step inference in Stable Diffusion: In the default leading timestep setting of DDIM, for single-step prediction, the model is provided with a timestep ($t\text{=}1$) suggesting the input is essentially denoised, yet in reality, the input is pure noise. The discrepancy between the timestep and the actual noise level causes the single-step result to be flawed. To align the timestep with the noise level received by the model, we adjust the scheduler to use the trailing setting ($t\text{=}T$ for single-step), correcting this issue and enabling single-step inference, significantly reducing the time overhead.
@@
Figure 3: Material Decomposition for 3D objects. Applying SuperMat’s powerful decomposition to 3D objects, we can obtain an object with high-quality materials in just three inference steps.
@@
$\begin{split}\mathcal{L}_{p}(\hat{y},y)=\sum_{j}\frac{1}{C_{j}H_{j}W_{j}}||\phi_{j}(\hat{y})-\phi_{j}(y)||_{1},\\ \mathcal{L}_{m}=\mathcal{L}_{p}(\hat{k}_{d},k_{d})+\mathcal{L}_{p}(\hat{k}_{rm},k_{rm}),\end{split}$ (2)
where $\phi$ is a 16-layer VGG network, and $\phi_{j}$ represents the output of the $j$-th layer, with a shape of $C_{j}\times H_{j}\times W_{j}$. In this paper, we use the $4$th, $8$th, $12$th, and $16$th layers to compute the feature reconstruction loss.
@@
Re-render Loss in DM Fine-tuning. Data-driven decomposition diffusion models typically focus on the distribution of individual material maps, yet they often overlook the interactions between different types of material maps. In other words, previous methods compute the loss between the predicted values and the ground truth without utilizing material information for further operations or calculations, which could provide stronger guidance for the model. As a result, although the predicted materials may appear close to the ground truth, the rendered output may not align with expectations. While this can be frustrating, it is understandable, as their training methods only produce meaningless predicted noise at each iteration. In contrast, our design, which incorporates structural expert branches and single-step prediction, allows us to obtain all denoised material maps during the training iterations. Using this information, we introduce the re-render loss, which involves rendering the model in a new lighting environment and comparing the result to the ground truth. Specifically, at each training iteration, we obtain $k_{d}$, $k_{rm}=(k_{r},k_{m})$, along with the true normal map $\hat{k}_{n}$, position map $\hat{k}_{p}$, camera viewpoint $\hat{c}$, and an environment light map. Then we use these information and the rendering model $\mathcal{R}$ described in Sec. 3.1 to compute the predicted material’s performance under new illumination. Similarly, with the groundtruth albedo $\hat{k}_{d}$ and RM $\hat{k}_{rm}$, we compute the ground truth of relighting. The perceptual loss between these two renderings defines our re-render loss:

### v4

@@ Abstract
Figure 1: We present an efficient method for high-quality material decomposition in both 2D and 3D. Left: Albedo and RM(Roughness, Metallic) predictions of SuperMat on rendered images under unknown illumination. The first row shows inputs with uneven lighting, the second row displays albedo predictions, and the third row presents RM predictions. Each texel in the RM map here is represented as $(1,\text{roughness},\text{metallic})$. Middle: Rendering results of a 3D object with PBR materials obtained by our method under novel lighting. Right: Our method is efficient and produces high-quality results.
@@ 1 Introduction
Decomposing Physically-Based Rendering (PBR) materials from images into their constituent properties (e.g., albedo, metallic, and roughness maps) remains a central challenge in computer vision and graphics. This task is essential for creating high-fidelity 3D assets across numerous applications, from film production and game development to virtual and augmented reality pipelines. While recent advances have shown promising results [6, 59, 30], the problem remains challenging due to its ill-posed nature, particularly given unknown lighting conditions and complex interactions between different material properties. Our work addresses two key objectives: 1) efficient decomposition of high-fidelity material maps from 2D images, and 2) extension of this capability to 3D objects, enabling accurate material estimation across complex geometries.
@@
Previous works in PBR material estimation generally fall into two main categories: optimization-based generation for 3D objects and data-driven decomposition for 2D images. Optimization-based methods [20, 23, 16, 28] employ differentiable rendering to reconstruct 3D objects with material parameters by minimizing the difference between rendered and target images. Instead, data-driven approaches learn the distribution of various materials from large datasets to perform intrinsic decomposition on 2D images, which is the primary focus of our work. Early methods  [49, 67] in this domain fail to produce stable and reliable results due to limited data availability and model capacity. Recent works  [26, 59, 6, 30] leverage large diffusion models and reframe material decomposition as a conditional image generation task. However, these methods tend to focus solely on individual material properties, overlooking the interactions between different types of materials and the combined impact on rendered appearances. Even minor inaccuracies in individual material estimates can propagate through the rendering pipeline, leading to substantial visual artifacts. Moreover, these diffusion-based methods face considerable computational overhead - they demand multiple denoising steps (30-50) during inference, causing a critical bottleneck for real-time applications.
@@
To address these challenges, we propose a deterministic one-step framework that achieves both efficient and high-quality material decomposition. First, we introduce and refine structural expert branches within the UNet architecture. These specialized branches enable concurrent processing of multiple material properties while maintaining domain-specific expertise, allowing our model to decompose albedo, metallic, and roughness maps simultaneously with minimal computational overhead. This design facilitates collaborative and physically consistent material prediction while avoiding the linear cost scaling typically required when estimating multiple material attributes. Second, we address the computational inefficiency of traditional diffusion models by reformulating the DDIM sampling process. Through modification of the scheduler, we achieve single-step inference without compromising quality, dramatically reducing inference time. Most importantly, this single-step formulation enables end-to-end training of the diffusion model, allowing us to compute the perceptual and re-render loss on final predicted material textures, which enhances shadow removal and improves the joint performance of decomposed materials under varying lighting conditions. These advances lead to SuperMat, a highly efficient image space material decomposition model that estimates high-fidelity albedo, metallic, and roughness maps at millisecond-scale speeds.
@@
For practical 3D applications, we further extend SuperMat’s capabilities to handle 3D objects. We upgrade SuperMat to a multi-view model, SuperMatMV, and design a UV refinement network that processes SuperMatMV’s predictions in UV space. This approach addresses the challenge of maintaining consistency and completeness across different, limited viewpoints while preserving the efficiency of our original 2D approach. Our pipeline processes a textured 3D model in approximately 3 seconds, producing high-quality material maps that maintain consistency across arbitrary viewpoints and lighting conditions.
@@
  • With the help of structural expert branches and single-step inference, we develop an end-to-end image space material decomposition training system that generates clean predictions for all material properties in each iteration, allowing for the integration of advanced techniques such as perceptual loss and re-render loss.
  • We propose an efficient PBR material decomposition method for 3D models. Combining SuperMatMV and our UV refinement network, this pipeline generates high-quality PBR materials for 3D objects from baked textures in approximately 3 seconds.
@@ 2 Related Works
Material Generation and Decomposition. Recognizing the importance of material information in 3D assets, researchers have increasingly focused on material acquisition for objects and scenes. On one hand, some 3D generation approaches [5, 52, 34, 39, 50, 55, 56, 65, 10] integrate material optimization into the appearance creation process, concentrating on generating PBR maps for untextured models. On the other hand, inverse rendering methods aim to reconstruct material-rich 3D representations from single or multi-view images. Some of these methods [60, 20, 7] assume a predefined lighting model to enhance estimation accuracy, but this constraint limits their practical applicability. To address this, other works [2, 63, 64, 21, 36, 23, 16, 62, 28] attempt to recover both material properties and unknown illumination simultaneously. As reliance on predefined assumptions decreases, researchers have found that beyond controlling material decomposition through rendered outputs, directly obtaining image space material priors is also feasible. Building on this idea, similar to ours, another body of research [49, 67, 26, 59, 6, 30, 32, 54, 46, 47, 45] focuses on material decomposition from 2D images captured or rendered under unknown lighting conditions. However, existing methods still face challenges related to training data, model architecture, and training strategies, leaving significant room for improvement in decomposition quality.
Diffusion Models. Recently, researchers have applied diffusion models to a range of tasks [38, 5, 48, 39, 65], demonstrating the effectiveness of casting outcome prediction as a conditional image generation process. By leveraging prior knowledge from large pre-trained models, recent studies [19, 43, 25, 12, 66, 15, 53, 6, 30] have achieved performance improvements over previous methods in tasks like depth, normal, and material estimation. However, incorporating large models introduces efficiency challenges, particularly high memory requirements and the time-intensive, multi-step denoising process needed for inference. Inspired by insights from recent work [31, 18, 29, 22], our approach addresses these issues by significantly reducing memory and inference time, improving the feasibility of applying these models in real-world applications.
Generation in UV Space. UV maps are advantageous for fully encapsulating the appearance information of 3D models, leading several works to explore 3D object manipulation directly within the UV space. By representing the texture of an object as a learnable image, these methods [17, 35, 41, 4, 58, 3] enable texture generation, inpainting, and editing, providing new avenues for 3D tasks and allowing 2D image-space techniques and insights to be applied to 3D appearances. Building on this approach, recent methods [1, 33, 58, 55, 57, 56, 13] have applied diffusion models to refine UV maps with promising results. In this spirit, to extend SuperMat’s capability to 3D, we fine-tune a UV-space refinement diffusion model specifically.
@@ 3 Method
In this paper, we propose SuperMat, an image space material decomposition model fine-tuned from Stable Diffusion that efficiently generates high-quality albedo, metallic, and roughness maps simultaneously. Given a rendered or captured image of an object under an unknown lighting condition from a specific viewpoint, SuperMat generates the corresponding material maps for the model from that same viewpoint. As shown in Figure 2, SuperMat is a 2D image-to-image one-step model with structural expert branches that simultaneously generates both albedo and RM maps, supporting single-step inference. Compared to previous diffusion-based methods, this model produces superior results in less time and with reduced memory usage. Our single-step formulation allows direct generation of material maps during training, in contrast to prior diffusion methods that predict noise ($\epsilon$) or velocity (v) terms. This enables end-to-end optimization with both perceptual loss on the generated maps and re-render loss from novel lighting conditions, leading to improved material quality and rendering fidelity. Building upon the single-view model, we further introduce a multi-view version, SuperMatMV, to ensure consistency in predictions across different viewpoints.
While SuperMat and SuperMatMV achieve efficient image space material decomposition, extending this capability to 3D objects presents additional challenges. Instead of relying on optimization-based approaches like score distillation sampling (SDS) that are computationally intensive and prone to mode collapse, we propose an efficient two-stage pipeline for material estimation on 3D objects. Given a 3D model with only RGB texture, we first apply SuperMatMV to decompose materials from multiple viewpoints. We then introduce a UV refinement network that consolidates these view-dependent predictions into complete, high-quality UV texture maps through single-step inference. As shown in Figure 3, this pipeline enables PBR material estimation for 3D objects in approximately 3 seconds per model.
@@ 3.1 SuperMat - Image Space Material Decomposition
Previous works [59, 6, 30] have demonstrated that casting material estimation as a conditional latent diffusion process, leveraging knowledge from large pre-trained diffusion models like Stable Diffusion [42], is effective. However, for our task that requires the simultaneous generation of multiple material maps, three key issues remain: 1) High model overhead: Fine-tuning Stable Diffusion with its original architecture can only estimate a single material type. Even when combining metallic and roughness into an RM map, treating it as a single material still requires two diffusion models, one for albedo and another for RM, to meet the rendering model’s requirements. This results in doubled training time and increased inference space overhead. 2) Slow inference speed: With the widely used DDIM scheduler, diffusion models require 30 to 50 denoising steps to produce satisfactory results, which can be inefficient and time-consuming. 3) Insufficient decomposition effect: Due to the high computational cost of backpropagation in multi-step inference, prior diffusion-based methods trained on predicted noise from single-step denoising rather than directly supervising the fully denoised material output. This limitation restricts the model’s ability to specialize in the decomposition task and impedes the application of more task-specific techniques.
@@
Structural Expert Branches. Our solution to the first problem is to introduce a structural expert branch architecture [31, 61]. Previous methods are limited to using a single diffusion model to estimate one material type. Although training separate models for different material maps is an intuitive approach, it introduces issues such as complex implementation, high overhead, and inconsistencies between materials. Considering that, despite their different representations, materials fundamentally reflect an understanding of object features, we incorporate structural expert branches into the UNet model, allowing a single diffusion model to predict multiple material maps. Specifically, we design two expert branches, one for predicting albedo, and the other for predicting RM. For each expert branch, we replicate the UNet’s last UpBlock, and conv_out, while the rest of the UNet architecture remains shared between the branches. The modified model first extracts and transforms encoded image features through the shared modules. These features are then separated into 2 copies by duplicating before the last UpBlock, with each branch ultimately predicting results for its respective domain. The expert branches specialize in material-specific features, while the shared modules focus on commonalities. In this way, we only add a small number of parameters (19,313,604 parameters, 2.23% of UNet’s total), achieving the same performance as previous methods that would otherwise require twice the scale.
@@
Figure 3: Material Decomposition for 3D objects. Applying SuperMat’s powerful decomposition to 3D objects, we can obtain an object with high-quality materials in just three inference steps. Due to implementation differences, the texels in the RM map for image space decomposition are $(1,\text{roughness},\text{metallic})$, while in the UV refinement process, they are $(0,\text{roughness},\text{metallic})$.
Single-Step Inference and End-to-End Training. To address the second problem, we analyze the DDIM scheduler of diffusion models. Previous works [18, 29, 51, 22] have highlighted a flaw in the implementation of the DDIM scheduler, which limits the potential of single-step inference in Stable Diffusion: In the default leading timestep setting of DDIM, for single-step prediction, the model is provided with a timestep ($t\text{=}1$) suggesting the input is essentially denoised, yet in reality, the input is pure noise. The discrepancy between the timestep and the actual noise level causes the single-step result to be flawed. To align the timestep with the noise level received by the model, we adjust the scheduler to use the trailing setting ($t\text{=}T$ for single-step), correcting this issue and enabling single-step inference, significantly reducing the time overhead.
@@
[…]
where $\phi$ is a 16-layer VGG network, and $\phi_{j}$ represents the output of the $j$-th layer, with a shape of $C_{j}\times H_{j}\times W_{j}$.
@@
Re-render Loss in DM Fine-tuning. Data-driven decomposition diffusion models typically focus on the distribution of individual material maps, yet they often overlook the interactions between different types of material maps. In other words, previous methods compute the loss between the predicted values and the ground truth without utilizing material information for further operations or calculations, which could provide stronger guidance for the model. As a result, although the predicted materials may appear close to the ground truth, the rendered output may not align with expectations. This limitation is inherent to their training methods, as they either produce meaningless predicted noise, preventing the application of image-level loss, or fail to generate all material predictions simultaneously. In contrast, our design, which incorporates structural expert branches and single-step prediction, allows us to obtain all denoised material maps during the training iterations. Using this information, we introduce the re-render loss [11], which involves rendering the model in a new lighting environment and comparing the result to the ground truth. Specifically, at each training iteration, we obtain $k_{d}$, $k_{rm}=(k_{r},k_{m})$, along with the true normal map $\hat{k}_{n}$, position map $\hat{k}_{p}$, camera viewpoint $\hat{c}$, and an environment light map. Then we use this information and our rendering model $\mathcal{R}$ to compute the predicted material’s performance under new illumination. Similarly, with the groundtruth albedo $\hat{k}_{d}$ and RM $\hat{k}_{rm}$, we compute the ground truth of relighting. The perceptual loss between these two renderings defines our re-render loss:
@@
Extension to Multi-view Material Decomposition. We extend SuperMat to a multi-view model to enhance the consistency of generated results across different viewpoints. Specifically, we use a 3D self-attention layer and incorporate camera extrinsics as conditions, following MVDream [44]. The multi-view model, SuperMatMV, is capable of simultaneously decomposing materials for 6 input views, providing more accurate and consistent material predictions.